# Consistent and Invariant Generalization Learning for Short-video Misinformation Detection

### v1

@@ Abstract.
Short-video misinformation detection has attracted wide attention in the multi-modal domain, aiming to accurately identify the misinformation in the video format accompanied by the corresponding audio. Despite significant advancements, current models in this field, trained on particular domains (source domains), often exhibit unsatisfactory performance on unseen domains (target domains) due to domain gaps. To effectively realize such domain generalization on the short-video misinformation detection task, we propose deep insights into the characteristics of different domains: (1) The detection on various domains may mainly rely on different modalities (i.e., mainly focusing on videos or audios). To enhance domain generalization, it is crucial to achieve optimal model performance on all modalities simultaneously. (2) For some domains focusing on cross-modal joint fraud, a comprehensive analysis relying on cross-modal fusion is necessary. However, domain biases located in each modality (especially in each frame of videos) will be accumulated in this fusion process, which may seriously damage the final identification of misinformation. To address these issues, we propose a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR), which contains two characteristic modules: (1) We involve the cross-modal feature interpolation to map multiple modalities into a shared space and the interpolation distillation to synchronize multi-modal learning; (2) We design the diffusion model to add noise to retain core features of multi modal and enhance domain invariant features through cross-modal guided denoising. Extensive experiments demonstrate the effectiveness of our proposed DOCTOR model. Our code is publicly available at https://github.com/ghh1125/DOCTOR.
@@
Hanghui Guo and Jia Zhu are with the Zhejiang Key Laboratory of Intelligent Education Technology and Application, Zhejiang Normal University
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandsubmissionid: 6136ccs: Information systems Multimedia information systemsccs: Computing methodologies Transfer learning
@@ 1. INTRODUCTION
Figure 2. The problem of Domain-bias Enhancement in Cross-modal Fusion. The red line is a simple case of bias accumulated (Audio bias and fusion bias accumulated are consistent with the video, so we have omitted them in this figure).
@@ 2.1. Multi-modal Misinformation Detection
Early research on multi-modal misinformation detection primarily focuses on text and image, emphasizing cross-modal correlation (Zhao et al., 2023; Zhang et al., 2023; Liu et al., 2024b; Zhang et al., 2024a; Wang et al., 2023; Ma et al., 2024). For example, MCAN (Wu et al., 2021) designs co-attention networks to fuse textual and visual features. CAFE (Chen et al., 2022) transforms features into a shared semantic space through cross-modal semantic alignment and combines cross-modal fuzzy learning to estimate the fuzziness between different modalities. MMICF (Zeng et al., 2023) introduces the causal inference module to eliminate the direct disparity in the fusion between text and image entities through causalities. Although these methods have shown promising performance in detecting misinformation involving text and images, their direct applicability to short-video misinformation is limited due to fundamental differences in content production mechanisms (Bu et al., 2023).
@@ 2.2. Multi-modal Domain Generalization
While uni-modal domain generalization has been well-studied (such as only text, image, or video), research on video-audio domain generalization remains limited (Ouali et al., 2023; Xu et al., 2022; Oh et al., 2023). Some of the recent progress on video-audio includes: SimMMDG (Dong et al., 2023) improves domain generalization by increasing the distance between modality-specific features and modality-shared features of audio and video. CMRF (Fan et al., 2024) improves domain generalization by reducing the flat minima between different modalities. RNA-Net (Planamente et al., 2022) balances audio and video feature norms via a relative norm alignment loss to enhance domain generalization. However, the existing domain generalization focus on video-audio methods is primarily tailored for general domains rather than specifically designed for misinformation detection, resulting in limited task adaptability and suboptimal performance.
@@ 3.1.1. Problem Formulation
Short-video misinformation detection. We assume that the input is a short video $v$ and accompanying audio $a$ (sometimes matched with a text description $t$ about the video title or introduction). After judging whether it contains misinformation, the output predicted label is represented as $Y\in\{0,1\}$, where 00 represents misinformation and $1$ represents the opposite. Building a misinformation detection model $\mathcal{M}$, input multi-modal data $\{v,a,t\}$, output predicted labels $\hat{Y}$, we minimize the loss function $\mathcal{L}(\cdot)$ between the predicted label $\hat{Y}$ and the true label $Y$ by optimizing the parameter $\Theta$. The following is the target formula:
@@ 3.2. Cross-modal Interpolation Distillation
(6) $\begin{split}L_{\text{dis}}^{v}=\left\|\hat{X}_{v}-\dot{X}_{v,a}\right\|_{2}^{2},\quad\theta\leq 0.5\\ L_{\text{dis}}^{a}=\left\|\hat{X}_{a}-\dot{X}_{v,a}\right\|_{2}^{2},\quad\theta>0.5\end{split}$
@@ 3.3.1. Multimodal Feature Fusion
For domain invariance of the spatial features in the video, we consider that creators of misinformation often add misleading elements, such as subtitles, to disguise forged frames. Thus, we select text-rich frames ($F_{d},d\in D$, where $D$ represents the set of all text-rich frames) as starting points. The extraction process for spatial features is detailed in Algorithm 1 (for the detailed steps, see Supplementary Material A). Following the steps below, we obtain the spatial features $X_{spa}$.
@@
0:  Frame $F_{d}$, Text Area ($TA$)
@@
2:  Apply CRAFT OCR to detect $TA$ in frame $F_{d}$
3:  Convert $TA$ to prompt embeddings $X_{vt}^{d}$ using SAM (Segment Anything Model) prompt encode
@@
[…]
@@
To further capture temporal dynamics, we introduce two encoding strategies: Temporal Position Encoding (TPE) and Duration Encoding (DE). TPE uses equal-frequency binning to assign learnable embeddings based on both absolute and relative segment durations. DE employs sinusoidal functions to represent segment positions within the video. These obtain $TPE^{i}$ and $DE^{i}$ of the segment $i$.
The final segment representation is constructed as:
(9) $Seg^{i}=DE^{i}+TPE^{i}+Seg_{t}^{i}+Seg_{v}^{i}+Seg_{a}^{i}.$
Lastly, all segments are then fused via self-attention (Self-Att) to obtain the overall temporal representation $X_{time}$:
@@ 4.1.1. Datasets
We use the Qwen2.5-72B (Yang et al., 2024) combined with manual verification for label classification. We divide into nine domains (Society, Health, Disaster, Culture, Education, Finance, Politics, Science, and Military). See Supplementary Material B for the detailed steps.
@@ 4.1.3. Implementations
All of our experiments were conducted on an NVIDIA RTX A6000. We set the learning rate as 5e-5 for FakeSV and 1e-3 for FakeTT. Specifically, for more information, see Supplementary Material C.
@@ 4.2.1. Domain generalization experiments
Specifically, compared with the recent state-of-the-art method for general domain generalization in video and audio, CMRF, our proposed DOCTOR model demonstrates a superior ability to effectively leverage multi-modal information while mitigating domain bias, thereby enhancing domain generalization in misinformation detection. As there is currently a lack of domain generalization methods tailored for short-video misinformation detection, we additionally compare the DOCTOR model with MMDFEND and MDFEND, which are SOTA methods (Open Source) for domain generalization in text-image misinformation detection. Across various domains, DOCTOR consistently achieves significantly higher accuracy.
@@
Finally, we conducted an ablation study on two key components of DOCTOR, Diffusion and Distillation, and found that the accuracy rate decreased. This indicates that these two components have an indispensable and effective synergy in the overall DOCTOR. They work together to provide key support for domain generalization.
Overall, DOCTOR outperforms other state-of-the-art methods in domain generalization for the following reasons: 1) DOCTOR designs interpolation knowledge distillation to solve the problem of modal competition caused by the need to make full use of multi-modal information when misinformation in different fields tends to be forged in different modalities. 2) DOCTOR is designed to diffuse comprehensive modal information, which effectively solves the problem of domain biases between different domains during domain collaboration. These designs fully improve the domain generalization ability of short-video misinformation detection.
@@ 4.2.2. General domain experiments
Another point worth discussing is whether solving the differences between different domains will interfere with misinformation detection in the general domain. As shown in Table 2, we conducted a thorough comparison using multiple indicators and multiple methods on two short-video misinformation detection datasets, FakeTT and FakeSV.
In the FakeSV dataset, DOCTOR achieved 84.91% accuracy and 84.70% F1. The overall performance is better than most existing short-video misinformation detection methods and some large language models. From the perspective of truth and misinformation, the DOCTOR is more effective in improving the method of identifying truth (Precision: 83.46%, Recall: 82.57%, F1: 83.67%). DOCTOR is also very competitive in identifying misinformation methods.
In the FakeTT dataset, DOCTOR also has great advantages, achieving 78.92% of accuracy and 78.04% of F1. This result surpasses other methods. Similarly, as mentioned above, there are similar performance trends in the real class and fake class.
@@ 4.2.3. Ablation experiments
Table 3. Ablation Study Results of DOCTOR model on FakeSV and FakeTT Datasets in the general domain.
@@
Table 4. Learning rate sensitivity results for DOCTOR model on FakeSV and FakeTT datasets (Accuracy).
Dataset 1e-5 5e-5 1e-4 5e-4 1e-3 5e-3
FakeSV 82.32 84.91 83.45 80.78 79.53 78.44
FakeTT 72.24 74.31 76.86 78.57 78.92 77.39
4.2.4. Parameter sensitivity analysis
@@ Acknowledgment
This work was supported by the National Key R&D Program of China under Grant No. 2022YFC3303600, the Zhejiang Provincial Natural Science Foundation of China under Grant No. LY23F020010, and the National Natural Science Foundation of China under Grant Nos. 62337001.
Appendix A Spatial Feature Extraction
We initially utilize a CRAFT OCR (Baek et al., 2019) to outline the text area ($TA$) of the frame $F_{d}$. Then we adopt the prompt encoder heuristic in the Segment Anything Model (SAM) (Kirillov et al., 2023) to convert $TA$ into prompt embeddings $X_{vt}^{d}$ to obtain the feature vector of the location and content information of the text area. At the same time, we input $F_{d}$ into the pre-trained ViT to generate the initial encoding $X_{v}^{d}$ to capture the spatial features of the image in a single frame.
For the obtained $X_{vt}^{d}$ and $X_{v}^{d}$, we optimize the representation of visual features $X_{v}^{d}$ through a bidirectional attention mechanism, making it more focused on the relative representation of subtitle features in space. The process mainly takes the following steps:
We input $X_{vt}^{d}$ and $X_{v}^{d}$ into a bidirectional attention block. The bidirectional attention block first performs a self-attention (Self-Att) operation on $X_{vt}^{d}$ to enhance its own feature representation.
(16) $\hat{X}_{vt}^{d}=\frac{(X_{vt}^{d}+\text{Self-Att}(X_{vt}^{d}))-\mu}{\sqrt{\sigma^{2}+\epsilon}}\sim\text{NORM}(X_{vt}^{d}+\text{Self-Att}(X_{vt}^{d}))$
where Self-Att is self-attention, and we normalize all features within the sample to have zero mean and unit variance to help alleviate the problem of internal covariate shift. Later, we use NORM to simplify this normalization process. Then, we perform cross-attention (Cross-Att) operations on the optimized features $\hat{X}_{vt}^{d}$ and $X_{v}^{d}$ to interact with each other and further optimize the $\hat{X}_{vt}^{d}$.
(17) $\dot{X}_{vt}^{d}=\text{NORM}(\hat{X}_{vt}^{d}+\text{Cross-Att}(\hat{X}_{vt}^{d},X_{v}^{d}))$
Lastly, a cross-attention operation is performed on $X_{v}^{d}$ to make it interact with the updated $\dot{X}_{vt}^{d}$ from MLP to obtain a more accurate visual feature representation.
(18) $\hat{X}_{v}^{d}=\text{NORM}(X_{v}^{d}+\text{Cross-Att}(\text{MLP}(\dot{X}_{vt}^{d}),X_{v}^{d}))$
After the attention process, the updated $\hat{X}_{v}^{d}$ is down-sampled through two convolutional layers and then expanded to obtain the spatial pattern features of a single frame $X^{d}$. Finally, we use self-attention to fuse these single-frame spatial features and obtain a unified spatial feature representation through average pooling (MEAN).
(19) $X_{spa}=\text{MEAN}(\text{Self-Att}([X^{1},X^{2},...,X^{d}])$
Appendix B Domain datasets
Table 5. Domains division of FakeSV and FakeTT datasets.
Domain FakeSV FakeTT
Culture 278 325
Disaster 1836 182
Education 195 28
Finance 116 9
Health 841 121
Politics 296 715
Military 121 84
Science 234 271
Society 1621 337
We use the Qwen2.5-72B (Yang et al., 2024) combined with manual verification for label classification based on the text and fakesv’s video collection introduction (Qi et al., 2023). We divide the above two datasets into nine domains, respectively. Specific domains’ names and numbers are detailed and shown in Table 5.
The specific prompt template is: “Here are some fields: Society, Health, Disaster, Culture, Education, Finance, Politics, Science, Military. And some examples from different domains [Some examples (Society: [case text], Health: [case text], …)]. You can then compare and analyze the [input text] with the examples above to choose the most suitable Domain.”
We divide the above two datasets into nine domains, respectively. Specific domains’ names and numbers are detailed and shown in Table 5. Since the number of some domains is small (Education, Finance, and Military in the FakeTT dataset), we conduct merging experiments with them.
Appendix C More Experimental Analyses
C.1. Training Strategy
We define the total loss function $Loss_{\text{final}}$ as the combination of the Cross-modal Interpolation Distillation loss $Loss_{\text{dis}}$ and the Cross-modal Invariance Fusion loss $Loss_{\text{es}}$ (emotion and sentiment features) and $Loss_{\text{st}}$ (spatial and temporal features). The specifically $Loss_{\text{final}}$ we define is as follows:
(20) $Loss_{\text{final}}=\alpha Loss_{\text{es}}+\beta Loss_{\text{st}}+\gamma Loss_{\text{dis}},$
where $\alpha$, $\beta$, and $\gamma$ denote the weighting factors.
C.2. Experimental Implementations
Our experiments include two aspects: domain generalization experiments and general domain experiments. The comparison methods selected in the domain generalization experiments are mainly the best performing and latest methods selected from general domain experiments for comparison. All our experiments were conducted on an NVIDIA RTX A6000 (48G). The $\alpha$ in the loss functions $Loss_{final}$ is 0.1, $\beta$ is 3, and $\gamma$ is 0.05. We choose the learning rates of 5e-5 for FakeSV and 1e-3 for FakeTT.
For more details on the DOCTOR model, our anonymous source code is available at https://anonymous.4open.science/r/DOCTOR-D1D2.
C.3. Comparison of Domain Generalization Star Map
Figure 6. Comparative analysis of six misinformation detection methods (FakingRecipe, SV-FEND, OpEvFake, CMRF*, MMDFND*, and DOCTOR) across two distinct datasets: FakeSV and FakeTT. * represents the domain generalization method.
As shown in Figure 6, the experiment present a comparative analysis of six misinformation detection methods (FakingRecipe, SV-FEND, OpEvFake, CMRF*, MMDFND*, and DOCTOR) across two distinct datasets: FakeSV and FakeTT. The radar visualization effectively illustrates each method’s performance across multiple content categories, including Disaster, Society, Health, Culture, Politics, Science, Education, Finance, and Military for FakeSV, with a combined Education/Finance/Military category for FakeTT. The charts reveal DOCTOR consistently outperforming other methods across most categories in both datasets, particularly excelling in Military (93.65%) for FakeSV and Culture (88.31%) for FakeTT.
C.4. More comparative accuracy analysis
Figure 7. Accuracy comparison of different methods on FakeSV and FakeTT datasets.
Figure 8. F1 comparison of different methods on FakeSV and FakeTT datasets.
The above two figures of Figure 7 and Figure 8 show the comparison of the Accuracy and F1-score of different methods on FakeSV and FakeTT datasets. It can be seen that the DOCTOR method achieved the best performance on both datasets.
Figure 9. Recall comparison of different methods on FakeSV and FakeTT datasets (Fake class).
Figure 10. Precision comparison of different methods on FakeSV and FakeTT datasets (Fake class).
Figure 11. F1 comparison of different methods on FakeSV and FakeTT datasets (Fake class).
The above three figures of Figure 11, Figure 10, and Figure 9 show the comparison of the Precision, Recall, and F1-score of different methods on the Fake category on the FakeSV and FakeTT datasets. It can be seen that the DOCTOR method performs outstandingly in Recall and F1-score on the Fake category.
Figure 12. Recall comparison of different methods on FakeSV and FakeTT datasets (Real class).
Figure 13. Precision comparison of different methods on FakeSV and FakeTT datasets (Real class).
Figure 14. F1 comparison of different methods on FakeSV and FakeTT datasets (Real class).
The above three figures of Figure 14, Figure 13, and Figure 12 show the comparison of the Precision, Recall, and F1-score of the Real category of different methods on the FakeSV and FakeTT datasets. From the results, the DOCTOR method performs best in Precision while maintaining a high level in Recall and F1-score, and the overall performance is stable.
C.5. Discussion on DOCTOR Model Efficiency
Despite the advanced features and complex mechanisms integrated into the DOCTOR model, as shown in Figure 15, our experiments reveal that the computational overhead does not experience a significant increase. The incorporation of cross-modal interpolation distillation and multi-modal invariance fusion, while enhancing the model’s generalization ability across different domains, does not lead to a substantial rise in computational cost. This indicates that DOCTOR is able to maintain efficient performance, even with the addition of multiple modalities and the complexity of noise-based feature refinement. The efficiency of DOCTOR is crucial for real-world applications, where both accuracy and computational efficiency are paramount.
Figure 15. Discussion on DOCTOR Model Efficiency.
C.6. Case Study
Figure 16. Case: Bill Gates Briefed CIA on Mind-Altering Vaccine in 2005.
As shown in Figure 16. Case: Bill Gates Briefed CIA on Mind-Altering Vaccine in 2005.
In this case, the DOCTOR model demonstrates strong domain generalization capabilities by correctly identifying the video as misinformation, while traditional methods like FakingRecipe misclassify it as truth. They struggle to adapt when encountering unseen distributions, variations in video styles, or novel linguistic manipulations. In contrast, DOCTOR incorporates multi-modal invariance fusion and cross-modal interpolation distillation, allowing it to generalize across different domains by capturing latent relationships between text, audio, and video. This enables it to detect cross-modal inconsistencies, making it more robust against out-of-domain misinformation scenarios.
Appendix D Temporal Position Encoding
Temporal Position Encoding (TPE) is designed to capture the coarse temporal location of each segment within the entire video. To achieve this, we divide the video into $B$ equal-frequency temporal bins and assign each segment to one of them based on its position in the sequence. Each bin is associated with a learnable embedding vector, which is used to represent the temporal context of segments assigned to that bin. This encoding helps the model understand whether a segment is located at the beginning, middle, or end of the video, which can be crucial for modeling temporal editing patterns.
Algorithm 2 Temporal Position Encoding (TPE)
0:  Number of segments $n$, number of bins $B$, embedding dimension $d$
0:  Temporal position embeddings $\{TPE^{1},\dots,TPE^{n}\}$
1:  $s\leftarrow\left\lfloor\frac{n}{B}\right\rfloor$
2:  Initialize embedding matrix $E\in\mathbb{R}^{B\times d}$
3:  for $i=1$ to $n$ do
4:     $b_{i}\leftarrow\min\left(\left\lfloor\frac{i-1}{s}\right\rfloor,B-1\right)$
5:     $TPE^{i}\leftarrow E[b_{i}]$
6:  end for
7:  return  $\{TPE^{1},\dots,TPE^{n}\}$
Appendix E Duration Encoding
Duration Encoding (DE) encodes the fine-grained temporal position of each segment using a fixed sinusoidal scheme, similar to the original positional encoding used in Transformers. This method ensures that each segment has a unique continuous representation based on its position index. For each segment $i$, and for each embedding dimension $j$, the encoding alternates between sine and cosine functions of different frequencies. This encoding provides precise temporal cues, enabling the model to capture temporal dependencies and alignment across modalities.
Algorithm 3 Duration Encoding (DE)
0:  Number of segments $n$, encoding dimension $d$
0:  Duration encodings $\{DE^{1},\dots,DE^{n}\}$
1:  Initialize $DE\in\mathbb{R}^{n\times d}$ with zeros
2:  for $i=0$ to $n-1$ do
3:     for $j=0$ to $d-1$ do
4:        $\theta\leftarrow\frac{i}{10000^{\frac{2\cdot\lfloor j/2\rfloor}{d}}}$
5:        if $j$ is even then
6:           $DE^{i+1}[j]\leftarrow\sin(\theta)$
7:        else
8:           $DE^{i+1}[j]\leftarrow\cos(\theta)$
9:        end if
10:     end for
11:  end for
12:  return  $\{DE^{1},\dots,DE^{n}\}$
Appendix F Predication
Algorithm 4 Prediction via Feature Diffusion and Distillation Fusion
1:  Input: Time feature $X_{\text{time}}$, Spatial feature $X_{\text{spa}}$,
         Semantic feature $X_{\text{sem}}$, Emotional feature $X_{\text{emo}}$,
         Video feature $X_{\text{video}}$, Audio feature $X_{\text{audio}}$
2:  Output: Final prediction score $\hat{Y}_{\text{final}}$
3:  {Step 1: Feature fusion and diffusion}
4:  $X_{\text{all}}\leftarrow\text{Concatenate}(X_{\text{time}},X_{\text{spa}},X_{\text{sem}},X_{\text{emo}})$
5:  $X_{\text{diffused}}\leftarrow\text{ApplyDiffusion}(X_{\text{all}})$
6:  {Step 2: Time and space prediction branch}
7:  $X_{\text{ts}}\leftarrow\text{Extract}(X_{\text{diffused}},[\text{time},\text{spa}])$
8:  $\hat{Y}_{\text{st}}\leftarrow\text{MLP}_{\text{TS}}(X_{\text{ts}})$
9:  {Step 3: Semantic and emotional prediction branch}
10:  $X_{\text{se}}\leftarrow\text{Extract}(X_{\text{diffused}},[\text{sem},\text{emo}])$
11:  $\hat{Y}_{\text{es}}\leftarrow\text{MLP}_{\text{SE}}(X_{\text{se}})$
12:  {Step 4: Video-Audio distillation output}
13:  $\hat{Y}_{\text{dis}}\leftarrow\text{Distill}(X_{\text{video}},X_{\text{audio}})$
14:  {Step 5: Late fusion to get final prediction}
15:  $\hat{Y}_{\text{fusion}}\leftarrow\hat{Y}_{\text{st}}\times\tanh(\hat{Y}_{\text{es}})$
16:  $\hat{Y}_{\text{final}}\leftarrow\hat{Y}_{\text{fusion}}\times\hat{Y}_{\text{dis}}$
17:  return $\hat{Y}_{\text{final}}$
To generate the final prediction, we design a multi-branch architecture that integrates temporal, spatial, semantic, and emotional features through a diffusion-enhanced fusion mechanism. As illustrated in Algorithm 4, the input features $X_{\text{time}}$, $X_{\text{spa}}$, $X_{\text{sem}}$, and $X_{\text{emo}}$ are first concatenated and passed through a diffusion process to enhance inter-feature dependencies and latent structure alignment. The diffused representation $X_{\text{diffused}}$ is then split into two branches:
[…]
In parallel, we distill information from both video and audio modalities using a specialized distillation module, resulting in $\hat{Y}_{\text{dis}}$.
For final prediction, we adopt a late fusion strategy where the time-spatial output is modulated by the nonlinear activation of the semantic-emotional output: $\hat{Y}{\text{fusion}}=\hat{Y}{\text{st}}\cdot\tanh(\hat{Y}{\text{es}})$. This fusion is then further modulated by the distilled multi-modal prediction: $\hat{Y}{\text{final}}=\hat{Y}{\text{fusion}}\cdot\hat{Y}{\text{dis}}$. This design effectively captures complementary information from all feature domains and ensures robustness through multi-modal consistency.
For specific code implementation, please refer to the code link.

### v2

@@ Abstract.
Short-video misinformation detection has attracted wide attention in the multi-modal domain, aiming to accurately identify the misinformation in the video format accompanied by the corresponding audio. Despite significant advancements, current models in this field, trained on particular domains (source domains), often exhibit unsatisfactory performance on unseen domains (target domains) due to domain gaps. To effectively realize such domain generalization on the short-video misinformation detection task, we propose deep insights into the characteristics of different domains: (1) The detection on various domains may mainly rely on different modalities (i.e., mainly focusing on videos or audios). To enhance domain generalization, it is crucial to achieve optimal model performance on all modalities simultaneously. (2) For some domains focusing on cross-modal joint fraud, a comprehensive analysis relying on cross-modal fusion is necessary. However, domain biases located in each modality (especially in each frame of videos) will be accumulated in this fusion process, which may seriously damage the final identification of misinformation. To address these issues, we propose a new DOmain generalization model via ConsisTency and invariance learning for shORt-video misinformation detection (named DOCTOR), which contains two characteristic modules: (1) We involve the cross-modal feature interpolation to map multiple modalities into a shared space and the interpolation distillation to synchronize multi-modal learning; (2) We design the diffusion model to add noise to retain core features of multi modal and enhance domain invariant features through cross-modal guided denoising. Extensive experiments demonstrate the effectiveness of our proposed DOCTOR model.  [^1] Our code is publicly available at https://github.com/ghh1125/DOCTOR
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755809isbn: 979-8-4007-2035-2/2025/10submissionid: 6136ccs: Information systems Multimedia information systems
@@ 1. INTRODUCTION
Figure 2. The problem of Domain-bias Enhancement in Cross-modal Fusion. The red line is a simple case of bias accumulated.
@@ 2.1. Multi-modal Misinformation Detection
Early research on multi-modal misinformation detection primarily focuses on text and image, emphasizing cross-modal correlation (Zhao et al., 2023; Zhang et al., 2023; Liu et al., 2024b; Zhang et al., 2024a; Wang et al., 2023; Ma et al., 2024). For example, MCAN (Wu et al., 2021) designs co-attention networks to fuse textual and visual features. CAFE (Chen et al., 2022) transforms features into a shared semantic space through cross-modal semantic alignment and combines cross-modal to estimate the fuzziness between different modalities. MMICF (Zeng et al., 2023) eliminates the direct disparity in the fusion between text and image entities through causalities. Although these methods have shown promising performance in detecting misinformation involving text and images, their direct applicability to short-video misinformation is limited due to fundamental differences in content mechanisms (Bu et al., 2023).
@@ 2.2. Multi-modal Domain Generalization
While uni-modal domain generalization has been well-studied (such as only text, image, or video), research on video-audio domain generalization remains limited (Ouali et al., 2023; Xu et al., 2022; Oh et al., 2023). Some of the recent progress on video-audio includes: SimMMDG (Dong et al., 2023) increases the distance between modality-specific features and modality-shared features of audio and video. CMRF (Fan et al., 2024) improves domain generalization by reducing the flat minima between different modalities. RNA-Net (Planamente et al., 2022) balances audio and video feature norms via a relative norm alignment loss. However, the existing domain generalization focus on video-audio methods is primarily tailored for general domains rather than specifically designed for misinformation detection, resulting in limited task adaptability and suboptimal performance.
@@ 3.1.1. Problem Formulation
Short-video misinformation detection. We assume that the input is a short video $v$ and accompanying audio $a$ (sometimes matched with a text description $t$ about the video title or introduction). After judging whether it contains misinformation, the output predicted label is represented as $Y\in\{0,1\}$, where 0 represents misinformation and $1$ represents the opposite. Building a misinformation detection model $\mathcal{M}$, input multi-modal data $\{v,a,t\}$, output predicted labels $\hat{Y}$, we minimize the loss function $\mathcal{L}(\cdot)$ between the predicted label $\hat{Y}$ and the true label $Y$ by optimizing the parameter $\Theta$. The following is the target formula:
@@ 3.2. Cross-modal Interpolation Distillation
(6) $\begin{split}L_{\text{dis}}^{v}=\left\|\hat{X}_{v}-\dot{X}_{a,v}\right\|_{2}^{2},\quad\theta\leq 0.5\\ L_{\text{dis}}^{a}=\left\|\hat{X}_{a}-\dot{X}_{a,v}\right\|_{2}^{2},\quad\theta>0.5\end{split}$
@@ 3.3.1. Multimodal Feature Fusion
For domain invariance of the spatial features in the video, we consider that creators of misinformation often add misleading elements, such as subtitles, to disguise forged frames. Thus, we select text-rich frames ($F_{d},d\in D$, where $D$ represents the set of all text-rich frames) as starting points. The extraction process for spatial features is detailed in Algorithm 1. Following the steps below, we obtain the spatial features $X_{spa}$.
@@
0: Frame $F_{d}$, Text Area (TATA)
@@
2: Apply CRAFT OCR to detect TATA in frame $F_{d}$
3: Convert TATA to prompt embeddings $X_{vt}^{d}$ using SAM (Segment Anything Model) prompt encode
@@
17: Fuse all frame features: $\text{Self-Att}([X^{1},X^{2},\ldots,X^{d}])$
@@
To further capture temporal dynamics, we introduce two encoding strategies: Temporal Position Encoding (TPE) and Duration Encoding (DE). TPE uses equal-frequency binning to assign learnable embeddings based on both absolute and relative segment durations. DE employs sinusoidal functions to represent segment positions within the video. These obtain $TPE^{i}$ and $DE^{i}$ of the segment $i$. The final segment representation is constructed as: $Seg^{i}(DE^{i}+TPE^{i}+Seg_{t}^{i}+Seg_{v}^{i}+Seg_{a}^{i}).$ Lastly, all segments are then fused via self-attention (Self-Att) to obtain the temporal representation $X_{time}$:
@@ 4.1.1. Datasets
We use Qwen2.5-72B (Yang et al., 2024) combined with manual verification for classification. We divide into nine domains (Society, Health, Disaster, Culture, Education, Finance, Politics, Science, and Military).
@@ 4.1.3. Implementations
All of our experiments were conducted on an NVIDIA RTX A6000. More information see code link.
@@ 4.2.1. Domain generalization experiments
Specifically, compared with the recent SOTA method for general domain generalization in video and audio, CMRF, our proposed DOCTOR model demonstrates a superior ability to effectively leverage multi-modal information while mitigating domain bias, thereby enhancing domain generalization in misinformation detection. As there is currently a lack of domain generalization methods tailored for short-video misinformation detection, we additionally compare the DOCTOR model with MMDFEND and MDFEND, which are SOTA methods (Open Source) for domain generalization in text-image misinformation detection. Across various domains, DOCTOR consistently achieves significantly higher accuracy.
@@
Finally, we conducted an ablation study on two key components of DOCTOR, Diffusion and Distillation, and found that the accuracy rate decreased. This indicates that these two components have an indispensable and effective synergy in the overall DOCTOR.
Table 3. Learning rate for DOCTOR model (Accuracy).
Dataset 1e-5 5e-5 1e-4 5e-4 1e-3 5e-3
FakeSV 82.32 84.91 83.45 80.78 79.53 78.44
FakeTT 72.24 74.31 76.86 78.57 78.92 77.39
@@ 4.2.2. General domain experiments
As shown in Table 2, another point worth discussing is whether solving the differences between different domains will interfere with misinformation detection in the general domain. (1) In the FakeSV dataset, DOCTOR achieved 84.91% accuracy and 84.70% F1. The overall performance is better than most existing short-video misinformation detection methods and some large language models. From the perspective of truth and misinformation, the DOCTOR is more effective in improving the method of identifying truth (Precision: 83.46%, Recall: 82.57%, F1: 83.67%). DOCTOR is also very competitive in identifying misinformation methods; (2) In the FakeTT dataset, DOCTOR also has great advantages, achieving 78.92% of accuracy and 78.04% of F1. This result surpasses other methods. Similarly, as mentioned above, there are similar performance trends in the real class and fake class.
4.2.3. Parameter sensitivity analysis
@@ 4.2.4. Ablation experiments
Table 4. Ablation Study of DOCTOR in the general domain.
@@
4.2.5. Visualization results
@@ Acknowledgment
This work was supported by the National Key R&D Program of China under Grant No. 2022YFC3303600, the Zhejiang Provincial Natural Science Foundation of China under Grant No. LY23F020010, and the National Natural Science Foundation of China under Grant No. 62337001.